# Graph Chirp Signal and Graph Fractional Vertex-Frequency Energy Distribution

### v1

@@ VIII Conclusion
In this paper, we formally defined graph chirp signals, a previously undefined class of signals in GSP. We proposed the GFED to enhance the analysis of these signals, offering a more expressive representation of graph signals in the vertex-frequency domain. We explored the marginal distribution properties of GFED, further enhancing its theoretical foundation and interpretability. We also introduced the GFGD as a flexible and generalized vertex-frequency distribution, along with the reduced interference GFED, which effectively suppresses cross-terms interference and improves signal clarity. Furthermore, we proposed a graph chirp signal detection method based on GFED domain filtering and performed several experiments on graph chirp signals, demonstrating its robustness in detecting graph chirp signals under noisy conditions. Moreover, according to the marginal distribution properties of GFED, the proposed method can be applied to real-world data, showcasing its effectiveness in denoising tasks. The results confirm the effectiveness of our framework in both signal detection and noise reduction tasks, highlighting its potential for broader applications in graph signal analysis.
Appendix A Proof of the filter transfer function matrix (31)
@@
$=$ $\mathbb{E}\left\{\sum_{n}\sum_{k}\left|E^{a}_{\mathbf{x}}(n,k)-\sum_{i}\widehat{E^{a}_{\mathbf{y}}}(i,k)\widehat{H}(i,k)U(n,i)\right|^{2}\right\}$
@@
$\mathbf{U}\left(\mathbb{E}\left\{\widehat{E^{a}_{\mathbf{y}}}\circ\overline{\widehat{E^{a}_{\mathbf{y}}}}\right\}\circ\widehat{\mathbf{H}}\right)=$ $\mathbf{U}\left(\widehat{E^{a}_{\mathbf{x}}}\circ\mathbb{E}\left\{\widehat{E^{a}_{\mathbf{y},b}}^{*}\right\}\right),$ (35)
@@
Appendix B Proof of the optimal filter (32)
@@
$=$ $\left|\widehat{E^{a}_{\mathbf{x}}}(l,k)\right|^{2}$
@@
$+\sigma^{2}\left|\sum_{i}x(i)\overline{U_{a}(i,k)}\thinspace\overline{U(i,l)}\right|^{2}$
@@
$+\sigma^{4}\left|\sum_{i}\left|U_{a}(i,k)\right|^{2}U(i,l)\right|^{2}$
$+\sigma^{4}\sum_{i}\left|U_{a}(i,k)\right|^{2}\left|U(i,l)\right|^{2}$
@@
where $\mathbb{E}\left\{\left|\widehat{E^{a}_{\mathbf{y}}}(l,k)\right|^{2}\right\}$ is given by (B).

### v2

@@ IX Conclusion
In this paper, we formally defined graph chirp signals, a previously undefined class of signals in GSP. We proposed the GFED to enhance the analysis of these signals, offering a more expressive representation of graph signals in the vertex–fractional-frequency domain. We explored the marginal distribution properties of GFED, further enhancing its theoretical foundation and interpretability. We also introduced the GFGD as a flexible and generalized vertex–fractional-frequency distribution, along with the reduced interference GFED, which effectively suppresses cross-terms interference and improves signal clarity. Furthermore, we proposed a graph chirp signal detection method based on GFED domain filtering and performed several experiments on graph chirp signals, demonstrating its robustness in detecting graph chirp signals under noisy conditions. Moreover, according to the marginal distribution properties of GFED, the proposed method can be applied to real-world data, showcasing its effectiveness in denoising tasks. The results confirm the effectiveness of our framework in both signal detection and noise reduction tasks, highlighting its potential for broader applications in graph signal analysis.
Appendix A Proofs of Theorems 1–4
Proof of Theorem 1: The GFRFT of the graph chirp signal $\mathbf{u}_{k}^{a}$ is
$\widehat{\left(\mathbf{u}_{k}^{a}\right)}_{b}=\mathbf{F}_{G}^{b}\left(\mathbf{F}_{G}^{a}\right)^{-1}\mathbf{e}_{k}=\mathbf{u}_{k}^{a-b}.$ (54)
When $a=b$, we have
$\widehat{\left(\mathbf{u}_{k}^{a}\right)}_{a}=\mathbf{u}_{k}^{0}=\mathbf{e}_{k}.$ (55)
Proof of Theorem 2: The $\ell_{2}$–norm of the graph chirp signal is
$\|\mathbf{u}_{k}^{a}\|_{2}=\mathbf{e}^{\rm T}_{k}\left(\left(\mathbf{F}_{G}^{a}\right)^{-1}\right)^{\rm H}\left(\mathbf{F}_{G}^{a}\right)^{-1}\mathbf{e}_{k}=\mathbf{e}^{\rm T}_{k}\mathbf{e}_{k}=1.$ (56)
Proof of Theorem 3: The inner product of the graph chirp signals with different graph initial frequencies is
$\langle\mathbf{u}_{k}^{a},\mathbf{u}_{l}^{a}\rangle=\mathbf{e}^{\rm T}_{k}\left(\left(\mathbf{F}_{G}^{a}\right)^{-1}\right)^{\rm H}\left(\mathbf{F}_{G}^{a}\right)^{-1}\mathbf{e}_{l}=\mathbf{e}^{\rm T}_{k}\mathbf{e}_{l}=\delta_{kl}.$ (57)
Proof of Theorem 4: According to the GFRFT–invariance of graph chirp signals, the GFED of $\mathbf{u}_{k_{0}}^{a_{0}}$ is given by
$E_{\mathbf{u}_{k_{0}}^{a_{0}}}^{a}(n,k)$ $=u_{k_{0}}^{a_{0}}(n)\,\overline{\widehat{\left(u_{k_{0}}^{a_{0}}\right)}_{a}(k)}\,\overline{u_{k}^{a}(n)}$
$=u_{k_{0}}^{a_{0}}(n)\,\overline{u_{k_{0}}^{a_{0}-a}(k)}\,\overline{u_{k}^{a}(n)}.$ (58)
In particular, when $a_{0}=a$, we have
$u_{k_{0}}^{a_{0}-a}(k)=e_{k_{0}}(k),$ (59)
and therefore, we arrive the required result (31).
Appendix B Proof of Theorem 5
@@
$\mathbf{U}\left(\mathbb{E}\left\{\widehat{E^{a}_{\mathbf{y}}}\circ\overline{\widehat{E^{a}_{\mathbf{y}}}}\right\}\circ\widehat{\mathbf{H}}\right)=$ $\mathbf{U}\left(\widehat{E^{a}_{\mathbf{x}}}\circ\mathbb{E}\left\{\overline{\widehat{E^{a}_{\mathbf{y}}}}\right\}\right),$ (62)
@@
Appendix C Proof of Theorem 6
@@
$=$ $\left|\widehat{E^{a}_{\mathbf{x}}}(l,k)\right|^{2}+\sigma^{2}\left|\sum_{i}x(i)\overline{U_{a}(i,k)}\thinspace\overline{U(i,l)}\right|^{2}$
@@
$+\sigma^{4}\left|\sum_{i}\left|U_{a}(i,k)\right|^{2}U(i,l)\right|^{2}+\sigma^{4}\sum_{i}\left|U_{a}(i,k)\right|^{2}\left|U(i,l)\right|^{2}$
@@
where $\mathbb{E}\left\{\left|\widehat{E^{a}_{\mathbf{y}}}(l,k)\right|^{2}\right\}$ is given by (C).